OpenStudy (anonymous):

Will fan and medal least common denominator x^4 / x^2x+1 -8 / x^2+5x+4

OpenStudy (anonymous):

\[\frac{ x ^{4} }{ x ^{2}+2x+1 } \frac{ -8 }{ x ^{2}+5x+4 }\]

OpenStudy (anonymous):

@DatTyGerTho

OpenStudy (anonymous):

I'm so confused

OpenStudy (anonymous):

yeah me too

OpenStudy (larseighner):

If you factor \(x^2 + 2x +1\) it should give you a big clue what to try on the other one

OpenStudy (anonymous):

i tried can you walk me through it

OpenStudy (larseighner):

Factor \(x^2 + 2x + 1\). then divide \(x^2+5x+4\) by one of the factors.

OpenStudy (anonymous):

can anyone walk me through this

OpenStudy (anonymous):

I'm just as confused as you are

OpenStudy (anonymous):

exactly this crap is hard

OpenStudy (anonymous):

@Abhisar

OpenStudy (mathstudent55):

Start by factoring both denominators.

OpenStudy (anonymous):

i did after that i get confused

OpenStudy (mathstudent55):

What did you get for the factorizations of the denominators?

OpenStudy (anonymous):

(x+1)(x-1) (x+2)(x-2)

OpenStudy (zehanz):

The first denominator can be factored as \((x+1)^2\), so it has two factors x+1. The second denominator can be factored as \((x+1)(x+4)\), so it has factors x+1 and x+4.

OpenStudy (mathstudent55):

\(\dfrac{ x ^{4} }{ x ^{2}+2x+1 } ~~~~~ \dfrac{ -8 }{ x ^{2}+5x+4 }\) \(\dfrac{ x ^{4} }{ (x +1)(x + 1) } ~~~~~ \dfrac{ -8 }{ (x +4)(x + 1) }\) Those are the correct factorizations.

OpenStudy (anonymous):

okay so the answer would be (x+1)^2 (x+4a)

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

without the a lol

OpenStudy (anonymous):

i have alot more if you dont mind helping with them

OpenStudy (zehanz):

Common denominators, have at least two factors x+1 and one factor x+4.

OpenStudy (zehanz):

I'll leave you in the able hands of @mathstudent55 :D

OpenStudy (mathstudent55):

\(\dfrac{ x ^{4} }{ (x +1)(x + 1) } ~~~~~ \dfrac{ -8 }{ (x +4)(x + 1) }\) \(\dfrac{ x ^{4} }{ (x +1)^2 } ~~~~~ \dfrac{ -8 }{ (x +4)(x + 1) }\) Once you get the prime facotrizations of the denominators, the LCD is the product of all common prime factors with larger exponent and and non-common prime factors

OpenStudy (anonymous):

ill tag one of you in the next one but the answer is (x+1)^2 (x+4)

OpenStudy (mathstudent55):

Let's look a the prime factors: Common factors: (x + 1)^2 and (x + 1): choose the larger exponent ---> (x + 1)^2 Non=common factor: (x + 4) ---> choose it too LCD: \((x + 1)^2(x + 4)\)

OpenStudy (anonymous):

okay im gonna tag you in the next one or do you care if i just post it on here